OpenStudy (anonymous):

x (x+2)^2 (x-1)^4 Have to find zeros, y-intercept and sketch end behavior. (help i got part of it)

OpenStudy (anonymous):

Zeros are 0; multiplicity of 1 -3 m= 3 2 m=2

OpenStudy (anonymous):

then for y intecept f(0)=0 then i dont know what to do

OpenStudy (whpalmer4):

The y-intercept is simply the value of the expression when x = 0, plug it in and calculate it...

OpenStudy (anonymous):

I got zero..

OpenStudy (whpalmer4):

oh, I misinterpreted what you wrote. how did you come up with a zero at x = -3 and x = 2?

OpenStudy (whpalmer4):

\[x(x+2)^2(x-1)^4 = 0\]\[x = 0\]\[(x+2) = 0\]\[(x-1) = 0\]\[x = 0, 1, -2\]are the zeros, no?

OpenStudy (whpalmer4):

End behavior is just what happens when you stick in a really big positive number and a really big negative number.

OpenStudy (anonymous):

Is the part where i get f(0)=0 right?

OpenStudy (anonymous):

Then how would I graph it though?..